Your firm has a cost of equity of 12 percent, a cost of preferred of 9 percent, and a pre-tax cost of debt of 8.5 percent. Your target capital structure is 35 percent equity, 15 percent preferred stock and 50 percent debt. The firm has a beta of 1.2 and a tax rate of 34 percent. What is your firm’s weighted average cost of capital?

Please Provide full steps (No excel sheet)
Answer #1

After-tax cost of debt=8.5*(1-tax rate)

=8.5*(1-0.34)=5.61%

WACC=Respective cost*Respective weight

=(12*0.35)+(9*0.15)+(5.61*0.5)

=8.355%